A pole-vaulter approaches the takeoff point at a speed of 9.00 m/s. Assuming that only this speed determines

the height to which he can rise, find the maximum height at which the vaulter can clear the bar. (4.13m)


1
Expert's answer
2022-05-24T13:56:29-0400
"h=\frac{v_0^2}{2g}=\frac{9.00^2}{2*9.81}=4.13\:\rm m"